OpenStudy (anonymous):

Can someone help me with this Algebra 1 question? I'll give a medal. Solve for x: |2x + 6| − 4 = 20

OpenStudy (anonymous):

\[\left| 2x+6 \right|=24\] remove the absolute value by putting + or negative sign \[2x+6=\pm 24\]

OpenStudy (anonymous):

x= 9, -15

OpenStudy (anonymous):

Ali can you tell me how you did it?

OpenStudy (anonymous):

right answers

OpenStudy (anonymous):

ill draw it for you.

OpenStudy (anonymous):

Catch me I'm not sure what that symbol is inbetween the equal sign and 24

OpenStudy (anonymous):

it is + or -

OpenStudy (anonymous):

as absolute value remove signs,so it could be + or negative

OpenStudy (anonymous):

Did you mean to put the + instead but miss clicked?

OpenStudy (anonymous):